You roll a dice twice and add the numbers.

Find the probability that you get a result less than 6 given that the tosses resulted in an odd number.
Answer #1

Baues' Theorem: P(A | B) = P(A & B) / P(B)

Number of cases that result in less than 6 and toss is odd = 6

[(1,2), (2,1), (1,4), (3,2), (2,3). (4,1)]

Number of cases in which result is odd = Number of cases with one number is odd and other number is even

= 6 x 3

= 18

P(less than 6 | tosses resulted in odd number) = Number of cases that result in less than 6 and tosses are odd / Number of cases in which result is odd)

= 6/18

= 1/3